Aisha’s high-protein diet contains 2.05 oz carbohydrate, 5.43 oz protein and 0.325 oz of fat. How many kcal of energy does Aisha derive from this diet? (Carbohydrate 4 kcal/g; Protein 4 kcal/g; Fat 9 kcal/g)
Answer #1


2.05 oz carbohydrate = 58.12 grams carbohydrate           ( 1 oz(ounce) = 28.35 grams)

amount of energy gained from carbohydrate = 58.12 g * 4 kcal/g = 232.5 kcal

5.43 oz protein = 153.94 grams protein

amount of energy gained from protein = 153.94 g * 4 kcal/g = 615.76 kcal

0.325 oz fat = 9.21 grams fat

amount of energy gained from fat = 9.21 g * 9 kcal/g = 82.89 kcal


Total energy gained by Aisha from this diet = 232.5+615.76+82.89

                                            = 931.15 kcal